Consider an open economy with a fixed price level in which investment, export, import and government spending are assumed to be autonomous. Consumption is directly and linearly related to disposable income while tax is directly and linearly related to national income.



a) Work out the multiplier with change in investment [8 marks]

b) If investment increases by N$30m, marginal propensity to consume is 0.7, tax rate is 0.3 and marginal propensity to import is 0.1, by how much will national income increases or decreases? [8 marks]
1
Expert's answer
2020-12-08T09:51:12-0500

a) Every time there is an injection of a new need in a circular flow there may be a repetitive effect. This is because the injection of more money leads to more spending, which creates more money, and so on. The result of the multiplication means an increase in the final income from any new method of spending.

The amount of recurrence depends on the indirect indoor decisions to be used, called the food-side bias (mpc), or savings, which is called the conservation side (mps). It is important to remember that when money is spent, this money becomes someone else's money, and so on. Most of the limits reflect the amount of extra money allocated for certain activities, such as spending by UK companies, household savings, and spending on exports. For example, if 80% of all revenue in a given time is spent on UK products, the minimum food trend would be 80/100, which is 0.8.

The following standard formula for calculating multiplier uses the size of the limits, as follows:

"\frac{1}{1}-mpc"

Therefore, if consumers use 0.8 and save 0.2 for every £ 1 of additional income, the multiplication will be:

"\frac{1}{1}-0.8=\frac{1}{0.2}=5"

Therefore, the multiplier is 5, which means that every £ 1 of new income generates £ 5 of additional income.

The effect of duplication in the open economy

As well as calculating the frequency of how much more money is spent, we can also measure the recurrence in terms of how much more money goes into savings, and other withdrawals. A full 'open' economy has all the sectors, therefore, threefold withdrawals - savings, taxes and imports This is reflected in the extra savings limits (mps) and more government revenue - foreign tax rate (mtr) and overseas value - power small input (mpm).

By combining all withdrawals we obtain the marginal retraction strength (mpw). The multiplication can now be calculated by the following equation:

"\frac{1}{1}-mpw"

Use 'repetition effect'

The concept of repetition can be applied to any situation where there is a new injection in the economy. Examples of such situations include:

When the government supports the construction of a new road

Where there is an increase in exports

Where there is a decrease in interest rates or tax rates, or when the exchange rate decreases.

Lower or 'rewind'

Withdrawal of revenue from the circular flow will result in a declining duplication effect. Therefore, whenever there is a growing recession, such as an increase in spending, import duties or taxes, there is a potential for recurring economic downturns.


b) Change in national income = Multiplier "\times" Change in investment

Change in investment = N"\$" 30m

Multiplier "=\frac{1}{[1-c(1-t)]+g}"

In the above c is marginal propensity to consume (mpcd), g is marginal propensity to import (mpm) and t is the tax rate

Multiplier "=\frac{1}{[1-0.7(1-0.3)]+0.1}=\frac{1}{[1-0.7(0.7)]+0.1}=\frac{1}{0.51+0.1}=\frac{1}{0.61}=1.64"

Change in national income "=1.64\times 30"

Change in national income "=49.18"

So, the national income increases by "N" "\$" "49.18"